OpenStudy (anonymous):

What is the answer to 5k-9 =2/3j solver for j

OpenStudy (aum):

To solve for j, isolate j: multiply both sides by 3 divide both sides by 2

OpenStudy (anonymous):

Thank you

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

When you ask a question like this one, you need to be careful that your fractions are clear to other readers. If you use the equation editor, it makes it clearer. For example, \[5k - 9 = \frac{ 2 }{ 3 }j\] or did you mean \[5k - 9 = \frac{ 2 }{ 3j }\] If you mean the first, you solve this by multiplying both sides of the equation by the "Inverse" of the fraction. Turn 2/3 upside down, and you get 3/2. Multiply both sides of the equation by 3/2. Remember to multiply both terms on the left by 3/2. \[\frac{ 3 }{ 2 } (5k - 9) = \frac{ 3 }{ 2 } * \frac{ 2 }{ 3 }j\] 3/2 and 2/3 cancel each other out. You now have isolated "j" by itself. Now all you have to do is finish the problem. BUT if you meant the other equation, \[5k - 9 = \frac{ 2 }{ 3j }\]then it would be much harder to solve, so I'm going to assume you meant the first. Do learn to use the equation editor. I just learned to use it earlier today.